a box contains ten balls of like shape and size. Three are red, two are white, and five are blue. The balls are numbered from 1 to 10. If the balls are drawn at random without replacement for each problem, find the following probabilities.
P(one red)=
P(two red)=
P(one white)=
P(one white, then one red)=
P(number 6)=
P(one ball with number > 5)=
P(one ball with number < 4)=
P(2 or 7)=
P(one ball with an even number)=
P(one ball even, then one ball odd)=
P(red ball, then white ball, then blue ball)=
P(1 and 2 and 3)=

Answer by stanbon(75887)

a box contains ten balls of like shape and size. Three are red, two are white,
and five are blue. The balls are numbered from 1 to 10. If the balls are drawn at random without replacement for each problem, find the following probabilities.
P(one red)= 3/10
---------------------------
P(two red)= 3C2/10C2 = 3/45 = 1/15
------------------------------
P(one white)= 2/10 = 1/5
---------------------------------
P(one white, then one red)= (2/10)(3/9) = (1/5)(1/3) = 1/15
-----------------------------------
P(number 6) = 1/10
-----------------------------------
P(one ball with number > 5)= 5/10 = 1/2
------------------------------------------
P(one ball with number < 4)= 3/10
-----------------------------------------
P(2 or 7)= 2/10 = 1/5
---------------------------------------
P(one ball with an even number)= 5/10 = 1/2
-----------------------------------------
P(one ball even, then one ball odd)=(5/10)(5/9) = 5/18
--------------------------------------------------------
P(red ball, then white ball, then blue ball)= (3/10)(2/9)(5/8) = 30/720 = 1/24
--------------------------------------------------------
P(1 and 2 and 3)= (1/10)(1/9)(1/8) = 1/720
============================================
Cheers,
Stan H.
====================